OpenStudy (anonymous):

4. During an elastic collision between two balls, the force-time plot has Gaussian profile. Provide the equations of velocities of both balls during the collision.

OpenStudy (anonymous):

use the law of conservation of momentum \[m_1u_1 + m_2u_2 = m_1v_1 + m_2v_2 \]

OpenStudy (anonymous):

during the collision, force is a gaussian function of time (given). let this function be f(t) At any instant of time, for both balls, \[m*dv/dt = f(t)\] let initial velocities be u1, u2 and masses m2, m2. let final velocites be v1, v2. then, conservation of momentum, \[m1*u1 + m2*u2 = m1*v1 + m2*v2\] and conservation of energy says, \[1/2m1*u1^2 + 1/2m2*u2^2 = 1/2m1*v1^2 + 1/2m2*v2^2\] so we can calculate the final velocites v1, v2 from above. What the question really asks is - What are the velocites during the collision. \[dv/dt = f(t)/m1\] integration from time = 0 to arbitrarty time (t) during the collsion. limits are at t = 0, v = u1 , at time = t, velocity = v \[\int\limits_{u1}^{v} dv = 1/m1*\int\limits_{0}^{t}f(t)dt\] Similarly for mass m2, \[\int\limits_{u2}^{v} dv = 1/m1*\int\limits_{0}^{t}f(t)dt\] These are the equations of velocities DURING the collision

OpenStudy (anonymous):

Thanks zaphodplaysitsafe

OpenStudy (anonymous):

you're welcome, kindly note that by mistake, for the second mass, right at the last, where i wrote the equation for velocity of 2nd mass, i divided by m1. ofcourse, it should be m2 :)